# Disorder-induced proximate quantum spin ice phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$

### v1

@@ Appendix D Modeling of magnetic diffuse-scattering data
Figure A4: (a) Goodness-of-fit map ($\chi^{2}$) as a function of the ring-exchange parameter $g$ (defined in Eq. 11) and the theoretical tuning parameter $\mu/g$ for the quantum spin ice model. The quantum limit corresponds to $\mu/g=0$, where the best-fit value is $g=0.075(1)$ K. (b) Goodness-of-fit map ($\chi^{2}$) versus the nearest- and next-nearest-neighbor exchange interactions $J_{1}^{\mathrm{ex}}$ and $J_{2}^{\mathrm{ex}}$ (defined in Eq. 13). The best-fit parameters are $J_{1}^{\mathrm{ex}}=0.25(1)$ K and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K. (c) Magnetization versus field measured at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5.0 mg, 6.3 mg, and 5.6 mg for magnetic fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. (d,e) Temperature dependence of powder-averaged dc susceptibility measured on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, shown alongside SPINTERACT Paddison (2023) simulations using the best-fit parameters from panel (b).
Figure A5: Illustration of the sign convention for the fitted interaction parameters from the classical model using the SPINTERACT program Paddison (2023). An antiferromagnetic nearest-neighbor interaction $J_{1}>0$ in the local frame (defined in Eq. 9) favors the canonical “two-in–two-out” spin configuration (red arrows) on each tetrahedron. A ferromagnetic next-nearest-neighbor interaction $J_{2}<0$ in the local frame favors a “loop” structure, where spins on three next-nearest-neighbor sites around a hexagonal plaquette align either all clockwise or all counterclockwise relative to the plaquette center.
To model the magnetic diffuse scattering observed in single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, we use the SPINTERACT software suite Paddison (2023). For quantum and classical models, we write the Hamiltonian in terms of unit Pauli matrices or unit-length Ising variables, respectively, corresponding to an effective spin-$1/2$ on each Pr3+ site with an effective $g$-factor $g_{\parallel}=5.17$ along the local $\langle 111\rangle$ direction Princep et al. (2013). The local coordinate frames $(\hat{\mathbf{a}}_{i},\hat{\mathbf{b}}_{i},\hat{\mathbf{e}}_{i})$ are defined as in Ref. Ross et al. (2011):
@@
where $d$ indexes a particular dataset with weight $W_{d}$, $I_{i}^{\text{data}}$ and $I_{i}^{\text{calc}}$ denote the measured and calculated intensities for the $i$-th data point, respectively, and $\sigma_{i}$ is the corresponding experimental uncertainty. The scale factor $s_{d}$ for each dataset is determined via the linear least-squares expression given above. For the quantum model, we adopt the theoretical results of Ref. Benton et al. (2012), where the goodness-of-fit $\chi^{2}$ is plotted against the ring-exchange parameter $g$ in the effective Hamiltonian:
@@
For the classical model, we employ Onsager reaction-field theory, incorporating long-range dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K, following the convention of Ref. Paddison (2023). The total interaction $J_{n}$ between $n$-th neighbor $\sigma_{i}^{z}$ terms in Eq. 1 thus includes both dipolar and exchange contributions,
@@
Further-neighbor interactions, in particular $J_{2}$ and $J_{3}$, have also been invoked to explain the magnetic correlations observed below the spin-freezing transition in $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ Yavors’kii et al. (2008); Samarakoon et al. (2022). The Hamiltonian, excluding disorder terms, is then:

### v2

@@ Appendix D Modeling of magnetic diffuse-scattering data
Figure A4: (a) Goodness-of-fit map ($\chi^{2}$) as a function of the ring-exchange parameter $g$ (defined in Eq. 11) and the theoretical tuning parameter $\mu/g$ for the quantum spin ice model. The quantum limit corresponds to $\mu/g=0$, where the best-fit value is $g=0.075(1)$ K. (b) Goodness-of-fit map ($\chi^{2}$) versus the nearest- and next-nearest-neighbor exchange interactions $J_{1}^{\mathrm{ex}}$ and $J_{2}^{\mathrm{ex}}$ (defined in Eq. 13). The best-fit parameters are $J_{1}^{\mathrm{ex}}=0.25(1)$ K and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K. (c) Magnetization versus field measured at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5.0 mg, 6.3 mg, and 5.6 mg for magnetic fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. (d,e) Temperature dependence of powder-averaged dc susceptibility measured on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, shown alongside SPINTERACT [41] simulations using the best-fit parameters from panel (b).
Figure A5: Illustration of the sign convention for the fitted interaction parameters from the classical model using the SPINTERACT program [41]. An antiferromagnetic nearest-neighbor interaction $J_{1}>0$ in the local frame (defined in Eq. 9) favors the canonical “two-in–two-out” spin configuration (red arrows) on each tetrahedron. A ferromagnetic next-nearest-neighbor interaction $J_{2}<0$ in the local frame favors a “loop” structure, where spins on three next-nearest-neighbor sites around a hexagonal plaquette align either all clockwise or all counterclockwise relative to the plaquette center.
To model the magnetic diffuse scattering observed in single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, we use the SPINTERACT software suite [41]. For quantum and classical models, we write the Hamiltonian in terms of unit Pauli matrices or unit-length Ising variables, respectively, corresponding to an effective spin-$1/2$ on each Pr3+ site with an effective $g$-factor $g_{\parallel}=5.17$ along the local $\langle 111\rangle$ direction [28]. The local coordinate frames $(\hat{\mathbf{a}}_{i},\hat{\mathbf{b}}_{i},\hat{\mathbf{e}}_{i})$ are defined as in Ref. [8]:
@@
where $d$ indexes a particular dataset with weight $W_{d}$, $I_{i}^{\text{data}}$ and $I_{i}^{\text{calc}}$ denote the measured and calculated intensities for the $i$-th data point, respectively, and $\sigma_{i}$ is the corresponding experimental uncertainty. The scale factor $s_{d}$ for each dataset is determined via the linear least-squares expression given above. For the quantum model, we adopt the theoretical results of Ref. [40], where the goodness-of-fit $\chi^{2}$ is plotted against the ring-exchange parameter $g$ in the effective Hamiltonian:
@@
For the classical model, we employ Onsager reaction-field theory, incorporating long-range dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K, following the convention of Ref. [41]. The total interaction $J_{n}$ between $n$-th neighbor $\sigma_{i}^{z}$ terms in Eq. 1 thus includes both dipolar and exchange contributions,
@@
Further-neighbor interactions, in particular $J_{2}$ and $J_{3}$, have also been invoked to explain the magnetic correlations observed below the spin-freezing transition in $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ [48, 49]. The Hamiltonian, excluding disorder terms, is then: